OpenStudy (anonymous):

Order the following numbers from least to greatest. 5 squared, -0.1, (-5/3) 0.7, 2 squared

OpenStudy (anonymous):

5 Square, -0.1, (-5/3), 0.7, 2 squared sorry i forgot the comma between (-5/3) and 0.7

OpenStudy (texaschic101):

5^2 = 25 -0.1 -5/3 = -1.66 0.7 2^2 = 4 least to greatest : - 5/3 , -0.1, 0.7, 2^2, 5^2